# The Role of Model Confidence on Bias Effects in Measured Uncertainties

### v1

@@ Abstract
With the growing adoption of Large Language Models (LLMs) for open-ended tasks, accurately assessing epistemic uncertainty, which reflects a model’s lack of knowledge, has become crucial to ensuring reliable outcomes. However, quantifying epistemic uncertainty in such tasks is challenging due to the presence of aleatoric uncertainty, which arises from multiple valid answers. While bias can introduce noise into epistemic uncertainty estimation, it may also reduce noise from aleatoric uncertainty. To investigate this trade-off, we conduct experiments on Visual Question Answering (VQA) tasks and find that mitigating prompt-introduced bias improves uncertainty quantification in GPT-4o. Building on prior work showing that LLMs tend to copy input information when model confidence is low, we further analyze how these prompt biases affect measured epistemic and aleatoric uncertainty across varying bias-free confidence levels with GPT-4o and Qwen2-VL. We find that all considered biases induce greater changes in both uncertainties when bias-free model confidence is lower. Moreover, lower bias-free model confidence leads to greater underestimation of epistemic uncertainty (i.e. overconfidence) due to bias, whereas it has no significant effect on the direction of changes in aleatoric uncertainty estimation. These distinct effects deepen our understanding of bias mitigation for uncertainty quantification and potentially inform the development of more advanced techniques.  [^1] https://github.com/XinyiLiu0227/Uncertainty_Quantification_Bias
The Role of Model Confidence on Bias Effects in Measured Uncertainties
@@ 1 Introduction
Figure 1: Uncertainty between valid answers (42 and 15) reflects aleatoric uncertainty, while uncertainty between 40 and 42, or between 16 and 15, reflects epistemic uncertainty due to the model’s lack of knowledge.
@@
Much of the existing literature leverages uncertainty to estimate a model’s confidence in its answers (Guo et al., 2017; Malinin and Gales, 2020). Model uncertainty can stem from aleatoric uncertainty, epistemic uncertainty, or both. Importantly, only epistemic uncertainty is indicative of the model’s confidence, as it captures the limitations of the underlying knowledge. In contrast, aleatoric uncertainty stems from the irreducible randomness of the true answer distribution and persists even if the model has perfect knowledge. As such, the true goal of “uncertainty quantification” is to quantify the epistemic uncertainty. When two predictions exhibit similar total uncertainty, the one driven by aleatoric uncertainty indicates a more knowledgeable and confident model than one dominated by epistemic uncertainty. Figure 1 illustrates this distinction through an example where the model is uncertain for different underlying reasons.
@@
Traditional uncertainty quantification methods typically estimate total uncertainty, as they often operate under the single-answer assumption, where aleatoric uncertainty is absent. Yet in real-world scenarios with multiple valid answers, distinguishing between the two becomes crucial.
@@
In settings where each question has only one valid answer and uncertainty is thus purely epistemic, it may be intuitive that the presence of bias, namely spurious features that models rely on without understanding the true semantic meanings, can lead to inaccurate uncertainty estimation based on biased generation probabilities. Therefore, mitigating bias can improve the effectiveness of uncertainty quantification based on generation probabilities (Jiang et al., 2023). However, the potential presence of aleatoric uncertainty introduces additional complexity. Bias may also reduce aleatoric uncertainty by concentrating probability mass on a single or smaller subset of valid answers. In such cases, bias may reduce the noise introduced by aleatoric uncertainty, potentially facilitating a clearer estimation of epistemic uncertainty.
@@
We investigate whether mitigating prompt-introduced biases can enhance uncertainty quantification with the presence of aleatoric uncertainty, using GPT-4o, one of the most advanced multimodal LLMs. These biases arise from arbitrary and unavoidable choices in spurious features that do not alter the underlying semantics when using a single prompt, such as phrasing, answer position, verbalizer assignment, and image shape (Wang et al., 2023; Liu et al., 2024; Gavrikov et al., 2024; Ye et al., 2024). Our results show that bias mitigation consistently enhance uncertainty quantification with the presence of aleatoric uncertainty, without requiring access to the internal model state. Specifically, removing text-based biases boosts AUROC (Hanley and McNeil, 1983; McDermott et al., 2024) by approximately 7%. Motivated by this, we further examine how bias affects epistemic and aleatoric uncertainty separately.
@@
Earlier research predominantly tackles the aleatoric uncertainty from different phrasings of the same semantic meaning, often by semantic equivalence calculations (Kuhn et al., 2023; Farquhar et al., 2024; Lin et al., 2023). Recent work (Ahdritz et al., 2024; Yadkori et al., 2024) has shifted focus towards more general scenarios, where multiple distinct semantic meanings are valid (Jiang et al., 2022; Jia et al., 2024; Barandas et al., 2024). These two studies find that models are more likely to copy information from prompts under high epistemic uncertainty than under high aleatoric uncertainty, which may be interpreted as a form of confirmation bias (Nickerson, 1998; Shi et al., 2024). Therefore, we hypothesize that the impact of the prompt-introduced biases examined in our earlier experiments on epistemic uncertainty amplifies with lower true model confidence, whereas its impact on aleatoric uncertainty remains relatively insensitive to confidence levels.
@@
Most multi-label Natural Language Processing datasets were introduced early and are now well-studied, allowing LLMs to achieve near-perfect performance with minimal uncertainty. We therefore construct visual-language datasets where LLM performance is not yet saturated, enabling analysis of both text-based and image-based prompt biases.
@@
As illustrated in Figure 2, greater overestimation of confidence in lower-confidence instances, a pattern observed in human behavior (Sulistyawati et al., 2011), can undermine the robustness of the ranking performance of the measured confidence. In extreme cases, such distortions may even reverse the correct ranking: when the true confidence in A exceeds that in B, the biased estimated confidence incorrectly favor B over A. Therefore, we further examine this directional change in uncertainty. We find that epistemic uncertainty is significantly more likely to be underestimated (i.e., overestimation of confidence) under bias when the model is genuinely less confident. In contrast, model confidence does not significantly affect the direction of aleatoric uncertainty shifts under bias.
The distinct effects of bias on epistemic and aleatoric uncertainty deepen our understanding of bias mitigation for uncertainty quantification and may guide the development of more advanced methods.
@@ 3 The role of Bias in Uncertainty Quantification
Ahdritz et al. (2024) and Yadkori et al. (2024) both found that LLMs are more likely to copy input information under high epistemic uncertainty but not high aleatoric uncertainty. Inspired by these findings, we further analyze how these prompt-introduced biases impact each type of uncertainty estimation, aiming to provide deeper insight.
@@ 3.1 Epistemic and Aleatoric Uncertainty
Building on the proven effectiveness of semantic equivalence in addressing phrasing variability, particularly the use of LLM-based Natural Language Inference (Farquhar et al., 2024), we focus on the challenge of multiple valid answers with distinct meanings. We adopt a multiple-choice format with two semantically distinct correct options and two incorrect ones. This setup ensures sufficient data points while providing a conceptual framework for our analysis without first resolving semantic equivalence. For generalizing uncertainty quantification from classification to open-ended generation, please refer to Appendix B of Jiang et al. (2023).
@@
[…]
@@
Aleatoric $\text{Aleatoric Entropy}=-\sum_{i\in\text{correct}}\frac{P(i)}{P(\text{correct})}\log\frac{P(i)}{P(\text{correct})}$ (3)
@@
Entropy=Epistemic $\text{Entropy}=\text{Epistemic Entropy}+P(\text{correct})\times\text{Aleatoric Entropy}$ (4)
@@ 3.2 Prompt-Introduced Biases
The orientation of images can influence the predictions of vision models, a phenomenon known as orientation bias (Henderson and Serences, 2021; Ye et al., 2024).
@@ 3.3 Uncertainty Quantification in the Presence of Aleatoric Uncertainty
Unlike the mutual information approach proposed by the recent work (Yadkori et al., 2024), which injects hints into prompts to measure copying behavior, our method operates in a smaller search space by directly targeting biases in default prompts, avoiding broader searches. Specifically, we address both text- and image-based biases unavoidably introduced by a single prompt, as identified in prior work (Wang et al., 2023; Liu et al., 2024; Gavrikov et al., 2024; Ye et al., 2024).
@@ 3.4 Bias Effects on Measured Uncertainties
To assess the impact of bias, we compare entropy values from single prompts to those averaged over multiple bias-shuffled prompts (see Figure 3). Specifically, we measure: (1) bias impact as the absolute change in epistemic and aleatoric entropy, and (2) bias-induced overconfidence as the decrease in entropy from the averaged distribution to the single prompt. While the averaged distribution across bias-shuffled prompts may not be entirely bias-free, it is relatively bias-reduced reference (Wang et al., 2023; Liu et al., 2024) and we refer to it as “bias-free” for convenience.
@@ 4 Experiments
We adopt the AUROC metric for uncertainty quantification, following prior studies (Band et al., 2022; Kuhn et al., 2023; Lin et al., 2023; Farquhar et al., 2024). AUROC is robust to class imbalance and effectively captures the ranking performance (McDermott et al., 2024).
@@
With OpenAI’s closed-source LLMs now providing top-20 token probabilities, we compute prediction probabilities across all options directly, rather than approximating via sampling (Farquhar et al., 2024). We approximate bias-free model confidence by summing correct options from averaging probabilities across bias-shuffled prompts. We also extend our experiments by approximating model inconfidence using bias-free epistemic entropy (higher entropy indicates lower confidence), presented in Appendix A.5.
@@ 5.1 Uncertainty Quantification Through Bias Mitigation
Rephrase-based Entropy (proposed) 10 0.7779 0.6442
Reorder-based Entropy (proposed) 10 0.7844 0.6299
Relabel-based Entropy (proposed) 10 0.7665 0.6406
Rephrase+Reorder+Relabel-based Entropy (proposed) 10*3 0.8123 0.6588
Resize-based Entropy (proposed) 10 0.7605 0.6219
Rotate-based Entropy (proposed) 10 0.7565 0.6204
Noise-based Entropy (proposed) 10 0.7535 0.6252
Resize+Rotate+Noise-based Entropy (proposed) 10*3 0.7699 0.6287
Table 2: This table presents the AUROC scores for epistemic uncertainty quantification with GPT-4o. While the Repetitive-based method shows minimal improvement, mitigation of any single bias consistently enhances performance on both datasets. Furthermore, combining methods targeting different biases further improves performance over individual methods.
@@
When model confidence (self-perception) aligns with its true knowledge, it serves as a good estimate of the probability of correctness. As shown in Equation (4), the model’s total uncertainty incorporates both epistemic uncertainty that indicates model confidence, and aleatoric uncertainty which does not. We use GPT-4o to evaluate the trade-off that bias mitigation introduces between these two types of uncertainty for uncertainty quantification.
@@
Table 4: Both GPT-4o and Qwen2-VL exhibit greater overconfidence in epistemic uncertainty estimation due to bias when their confidence is lower, demonstrated by the negative coefficients and statistically significant p-values. In contrast, model confidence has no significant effect on the direction of aleatoric entropy changes caused by bias, supported by mostly insignificant p-values and mixed coefficient directions. The coefficient ratio Epi./Ale. > 1 is bolded. (***p $\leq$ 0.001, **p $\leq$ 0.01, *p $\leq$ 0.05, ns=not significant p > 0.05)
@@ 5.2 Relationship Between Model Confidence and Bias Impact
Similar results are obtained using bias-free epistemic entropy as the approximated model inconfidence, as shown in Appendix A.5.
@@ 5.3 Relationship Between Model Confidence and Bias-Induced Overconfidence
While lower model confidence leads to greater bias-induced changes, the direction of change is crucial. Greater under-confidence (i.e. higher measured entropy) in lower bias-free confidence instances improves the robustness of estimated confidence ranking under estimation noise by amplifying the contrast between instances with low and high bias-free confidence. However, greater over-confidence in lower-confidence instances hurts the ranking performance of estimated confidence (see Figure 2).
@@
Lower model confidence is associated with greater underestimation of epistemic entropy (i.e., overconfidence) in the presence of bias.
@@
Model confidence has no significant effect on the direction of aleatoric entropy changes caused by bias.
@@
Using bias-free epistemic entropy to approximate model inconfidence yields similar results, as shown in Appendix A.5.

### v2

@@ Abstract
With the growing adoption of Large Language Models (LLMs) for open-ended tasks, accurately assessing epistemic uncertainty, which reflects a model’s lack of knowledge, has become crucial to ensuring reliable outcomes. However, quantifying epistemic uncertainty in such tasks is challenging due to the presence of aleatoric uncertainty, which arises from multiple valid answers. While bias can introduce noise into epistemic uncertainty estimation, it may also reduce noise from aleatoric uncertainty. To investigate this trade-off, we conduct experiments on Visual Question Answering (VQA) tasks and find that mitigating prompt-introduced bias improves uncertainty quantification in GPT-4o. Building on prior work showing that LLMs tend to copy input information when model confidence is low, we further analyze how these prompt biases affect measured epistemic and aleatoric uncertainty across varying bias-free confidence levels with GPT-4o and Qwen2-VL. We find that all considered biases have greater effects in both uncertainties when bias-free model confidence is lower. Moreover, lower bias-free model confidence is associated with greater bias-induced underestimation of epistemic uncertainty, resulting in overconfident estimates, whereas it has no significant effect on the direction of bias effect in aleatoric uncertainty estimation. These distinct effects deepen our understanding of bias mitigation for uncertainty quantification and potentially inform the development of more advanced techniques.  [^1] https://github.com/XinyiLiu0227/Uncertainty_Quantification_Bias
The Role of Model Confidence on Bias Effects in Measured Uncertainties for Vision-Language Models
@@ 1 Introduction
Figure 1: Uncertainty between valid answers (e.g., France and Paris) reflects aleatoric uncertainty, while uncertainty between Paris and Tokyo reflects epistemic uncertainty due to the model’s lack of knowledge.
@@
Much of the existing literature leverages uncertainty to approximate a model’s confidence in its answers (Guo et al., 2017; Malinin and Gales, 2020). Model uncertainty can stem from aleatoric uncertainty, epistemic uncertainty, or both. Importantly, only epistemic uncertainty is indicative of the model’s confidence, since it captures the limitations of the model’s knowledge. In contrast, aleatoric uncertainty stems from the irreducible randomness of the true answer distribution and persists even if the model has perfect knowledge. As such, the true goal of “uncertainty quantification” is to quantify the epistemic uncertainty. When two predictions exhibit similar total uncertainty, the one driven by aleatoric uncertainty indicates a more knowledgeable and confident model than one dominated by epistemic uncertainty. Figure 1 illustrates this distinction through an example where the model is uncertain for different underlying reasons.
@@
Traditional uncertainty quantification methods, however, typically estimate total uncertainty. This is because they often operate under the single-answer assumption, where aleatoric uncertainty is absent. Yet in real-world scenarios with multiple valid answers, distinguishing between the two becomes crucial.
@@
In settings where each question has only one valid answer and uncertainty is thus purely epistemic, it may be intuitive that the presence of bias (Ye et al., 2024a; Yang et al., 2024; Seo et al., 2022), namely spurious features that models rely on without understanding the true semantic meanings, can lead to inaccurate uncertainty estimation based on biased generation probabilities. Therefore, mitigating bias can improve the effectiveness of uncertainty quantification based on generation probabilities (Jiang et al., 2023). However, the potential presence of aleatoric uncertainty introduces additional complexity. Bias may also reduce aleatoric uncertainty by concentrating probability mass on a single or smaller subset of valid answers. In such cases, bias may reduce the noise introduced by aleatoric uncertainty, potentially facilitating a clearer estimation of epistemic uncertainty.
@@
We investigate whether mitigating prompt-introduced biases can enhance uncertainty quantification with the presence of aleatoric uncertainty, using GPT-4o, one of the most advanced multimodal LLMs. These biases arise from arbitrary and unavoidable choices in spurious features that do not alter the underlying semantics when using a single prompt, such as phrasing, answer position, verbalizer assignment, and image shape (Wang et al., 2023; Liu et al., 2024; Gavrikov et al., 2024; Ye et al., 2024b). Our results show that bias mitigation consistently enhances uncertainty quantification with the presence of aleatoric uncertainty, without requiring access to the internal model state. Specifically, removing text-based biases boosts AUROC (Hanley and McNeil, 1983; McDermott et al., 2024) by approximately 7%. Motivated by this, we further examine how bias affects epistemic and aleatoric uncertainty separately.
@@
Earlier research predominantly tackles the aleatoric uncertainty from different phrasings of the same semantic meaning, often by semantic equivalence calculations (Kuhn et al., 2023; Farquhar et al., 2024; Lin et al., 2023). Recent work (Ahdritz et al., 2024; Yadkori et al., 2024) has shifted focus towards more general scenarios, where multiple distinct semantic meanings are valid (Jiang et al., 2022; Jia et al., 2024; Barandas et al., 2024). These two studies find that models are more likely to copy information from prompts under high epistemic uncertainty than under high aleatoric uncertainty, which may be interpreted as a form of confirmation bias (Nickerson, 1998; Shi et al., 2024). Therefore, we hypothesize that the impact of the prompt-introduced biases examined in our earlier experiments on epistemic uncertainty amplifies with lower bias-free model confidence, whereas its impact on aleatoric uncertainty remains relatively insensitive to confidence levels.
@@
Most multi-label Natural Language Processing datasets were introduced early and are now well studied, allowing LLMs to achieve near-perfect performance with minimal uncertainty (Yadkori et al., 2024). We therefore construct visual-language datasets where LLM performance is not yet saturated, enabling analysis of both text-based and image-based prompt biases.
@@
As illustrated in Figure 2, assigning more inflated confidence scores to lower-confidence instances undermines the robustness of ranking by measured confidence. Notably, this pattern has been observed in human behavior (Sulistyawati et al., 2011). In extreme cases, such distortions may even reverse the correct ranking: when the bias-free confidence in A exceeds that in B, the biased estimates incorrectly favor B over A. To better understand this directional distortion in confidence estimation, we examine how bias interacts with different sources of uncertainty. We find that bias-induced underestimation of epistemic uncertainty is greater when the model is genuinely less confident, resulting in overconfident estimates. In contrast, the bias-induced directional shift in aleatoric uncertainty estimation shows no significant association with confidence.
The distinct effects of bias on epistemic and aleatoric uncertainty deepen our understanding of bias mitigation in uncertainty quantification. This understanding may also guide the development of more advanced methods.
@@ 3 The Role of Bias in Uncertainty Quantification
Ahdritz et al. (2024) and Yadkori et al. (2024) both found that LLMs are more likely to copy input information under high epistemic uncertainty but not high aleatoric uncertainty. Inspired by these findings, we further analyze how these prompt-introduced biases impact each type of uncertainty estimation separately, aiming to provide deeper insight.
@@ 3.1 Epistemic and Aleatoric Uncertainty
Building on the proven effectiveness of semantic equivalence in addressing phrasing variability, particularly the use of LLM-based Natural Language Inference (Farquhar et al., 2024), we focus on the challenge of multiple valid answers with distinct meanings. To address this, we adopt a multiple-choice format with two semantically distinct correct options and two incorrect ones. This design provides sufficient data for analysis while offering a clear conceptual framework, without introducing additional variance from semantic-equivalence resolution. For generalizing uncertainty quantification from classification to open-ended generation, please refer to Appendix B of Jiang et al. (2023).
@@
$\text{Epistemic Entropy}=$ $-P(\text{correct})\log P(\text{correct})$
@@
$\text{Aleatoric Entropy}=-\sum_{i\in\text{correct}}\frac{P(i)}{P(\text{correct})}\log\frac{P(i)}{P(\text{correct})}$ (3)
@@
[…]
@@ 3.2 Prompt-Introduced Biases
The orientation of images can influence the predictions of vision models, a phenomenon known as orientation bias (Henderson and Serences, 2021; Ye et al., 2024b).
@@ 3.3 Uncertainty Quantification in the Presence of Aleatoric Uncertainty
Unlike the mutual information approach proposed by the recent work (Yadkori et al., 2024), which injects hints into prompts to measure copying behavior, our method operates in a smaller search space by directly targeting biases in default prompts, avoiding broader searches. Specifically, we address both text- and image-based biases unavoidably introduced by a single prompt, as identified in prior work (Wang et al., 2023; Liu et al., 2024; Gavrikov et al., 2024; Ye et al., 2024b).
@@ 3.4 Bias Effects on Measured Uncertainties
To assess the impact of bias, we compare entropy values from single prompts to those averaged over multiple bias-shuffled prompts (see Figure 3). Specifically, we measure: (1) bias effect as the absolute change in epistemic and aleatoric entropy, and (2) bias-induced underestimation as the decrease in entropy from the averaged distribution to the single prompt. While the averaged distribution across bias-shuffled prompts may not be entirely bias-free, it is relatively bias-reduced reference (Wang et al., 2023; Liu et al., 2024) and we refer to it as “bias-free” for convenience.
@@ 4 Experiments
We adopt the AUROC metric for uncertainty quantification, following prior studies (Band et al., 2022; Kuhn et al., 2023; Lin et al., 2023; Farquhar et al., 2024). AUROC measures how well confidence scores rank instances with correct versus incorrect predictions and is robust to class imbalance (McDermott et al., 2024).
@@
Due to budget constraints, we evaluate only two randomly selected biases for Claude-3.7-Sonnet (‘claude-3-5-sonnet-20241022’) on CREPE, with the results presented in Appendix A.5.
@@
With OpenAI’s closed-source LLMs now providing top-20 token probabilities, we compute prediction probabilities across all options directly for GPT-4o. When token probabilities are not exposed, a common strategy is to approximate them through sampling, which has been shown effective (Farquhar et al., 2024). Following this approach, we approximate the token probabilities for Claude-3.7-Sonnet using 10 sampled prompts.
Bias-free model confidence is estimated by summing the probabilities of correct options averaged across bias-shuffled prompts. We further extend our experiments by approximating model inconfidence using bias-free epistemic entropy (higher entropy indicates lower confidence), with the results presented in Appendix A.5.
@@ 5.1 Uncertainty Quantification Through Bias Mitigation
[…]
Table 2: This table presents the AUROC scores for uncertainty quantification with GPT-4o. While the Repetitive-based method shows minimal improvement, mitigation of any single bias consistently enhances performance on both datasets. Furthermore, combining methods targeting different biases further improves performance over individual methods.
@@
When model confidence (self-perception) aligns with its true knowledge (Kadavath et al., 2022; Farquhar et al., 2024), it serves as a good estimate of the probability of correctness. As shown in Equation (4), the model’s total uncertainty incorporates both epistemic uncertainty that indicates model confidence, and aleatoric uncertainty which does not. We use GPT-4o to evaluate the trade-off that bias mitigation introduces between these two types of uncertainty for uncertainty quantification.
@@
Table 4: Both GPT-4o and Qwen2-VL exhibit greater bias-induced underestimation of epistemic uncertainty estimation when their confidence is lower, demonstrated by the negative coefficients and statistically significant p-values. In contrast, model confidence has no significant effect on the direction of bias effect in aleatoric entropy, supported by mostly insignificant p-values and mixed coefficient signs. The coefficient ratio Epi./Ale. > 1 is bolded. (***p $\leq$ 0.001, **p $\leq$ 0.01, *p $\leq$ 0.05, ns=not significant p > 0.05)
@@ 5.2 Relationship Between Model Confidence and Bias Impact
Similar results are obtained using bias-free epistemic entropy as the approximated model inconfidence, as shown in Appendix A.5. The extended experiments on Claude-3.7-Sonnet with two randomly selected biases also align with these findings and are presented in the same appendix.
@@ 5.3 Relationship Between Model Confidence and Bias-Induced Overconfidence
While lower model confidence leads to greater bias-induced changes, the direction of change is crucial. Greater under-confidence (i.e. overestimation of uncertainty) in lower bias-free confidence instances improves the robustness of estimated confidence ranking under estimation noise by amplifying the contrast between instances with low and high bias-free confidence. However, greater over-confidence in lower-confidence instances hurts the ranking performance of estimated confidence (see Figure 2).
@@
Lower model confidence is associated with greater bias-induced underestimation of epistemic entropy (i.e., overconfidence).
@@
Model confidence has no significant effect on the direction of bias-induced aleatoric entropy changes.
@@
Using bias-free epistemic entropy to approximate model inconfidence yields similar results, as shown in Appendix A.5. The extended experiments on Claude-3.7-Sonnet with two randomly selected biases also align with these findings and are presented in the same appendix.